# Less but Better: Enabling Generalized Zero-shot Learning Towards Unseen Domains by Intrinsic Learning from Redundant LLM Semantics

### v1

@@ I Introduction
The era of Large Language Models (LLMs) marks a tremendous leap in the ability of artificial intelligence to understand and generate natural language. With massive numbers of parameters and extensive training data (for example, GPT-3 has 175 billion parameters and 753.4 GB of training data [1]), LLMs have the capacity to internalize vast amounts of knowledge and externalize it for application in the real world [2]. In some machine learning studies that fuse data and knowledge [3, 4, 5], researchers previously relied on the knowledge from human experts to provide models with prior information, aiming to obtain results that outperform purely data-driven techniques. The emergence of LLMs has, to some extent, supplanted the role of human experts, thereby bringing promising prospects to data-knowledge fusion studies, such as zero-shot learning (ZSL) [6, 7, 8, 9].
ZSL is designed to recognize unseen classes not present in the training set by leveraging semantic knowledge to bridge the gap between seen and unseen classes, mimicking human cognitive inference [10, 11, 12]. Initially, ZSL approaches like DAP [11] and ALE [13] construct mappings from data to semantics, training on seen classes and extrapolating to unseen test data, under the assumption that only unseen classes would be tested. This works well when the assumption holds, but real-world scenarios often involve a mix of seen and unseen classes. To overcome this limitation, Generalized Zero-Shot Learning (GZSL) [14, 8, 15, 16] is introduced, requiring models to identify both seen and unseen classes during testing. However, ZSL models tend to misclassify unseen classes as seen due to the domain shift problem (DSP) [6, 17], leading to poor GZSL performance. To combat this, generative methods such as f-CLSWGAN [18] and f-vaegan-d2 [19] have been developed, modeling between the semantic space and the feature space, generating unseen class data based on semantic knowledge to reduce model overfitting on seen classes and mitigate DSP.
@@
ZSL
Single
Unseen class only
GZSL
Single
Seen & unseen class
CDZSL
Multiple
Unseen class only
CDGZSL
Multiple
Seen & unseen class
@@ III-F The CDGZSL inference procedure
Because $C_{c}$ is trained by both real seen features and fake unseen features with less effection of semantic variation, the DSP could be further alleviated, and we could gain a more accurate classifier of CDGZSL.
@@ IV-A2 Mini-DomainNet
Semantics Methods Art Clipart Product
@@
LLM text CuMix
img-only
59.43 0 0 50.91 3.28 6.16 66.21 0.72 1.42
two-level
57.37 1.85 3.58 50.77 3.61 6.74 66.88 0.70 1.38
original
57.83 2.46 4.71 51.19 3.94 7.31 65.09 1.80 3.50
@@
Attribute CuMix
img-only
43.81 4.93 8.86 41.10 2.96 5.52 56.49 2.52 4.82
two-level
43.45 5.56 9.85 39.35 4.60 8.23 55.17 6.49 11.61
original
44.28 6.17 10.83 38.95 7.23 12.19 53.16 9.38 15.94
@@ IV-B Experimental details
In CDGZSL, there are both seen and unseen domains, as well as seen and unseen categories. For Office-Home, the allocation of seen and unseen domains is as follows: Art, Clipart, and Product are each designated as the unseen domain in turn. When a domain is designated as the unseen domain, the other three domains serve as the seen domains. Because we employ Resnet-50[34] pretrained on Imagenet as the initial parameters for the data encoder $E_{d}$, only the Real-world domain is not used as the unseen domain (since it is more similar to Imagenet) among these four domains. Additionally, the division of seen and unseen categories is as follows: we select 5 classes that do not appear in Imagenet as unseen classes, 5 classes as validation, and the remaining 55 classes as seen classes. For Mini-DomainNet, similarly, aside from the Real domain, the other three domains each take a turn as the unseen domain once. Also, we select 13 classes that do not appear in Imagenet as unseen classes, 13 classes as validation, and the remaining classes as seen classes.
For the various networks used in the experiments, the specific details are as follows: The data encoder $E_{d}$ is the encoder of Resnet50, with initial parameters trained on Imagenet; The domain classifier $C_{d}$, seen class classifier $C_{s}$, and the final CDGZSL classifier $C_{c}$ are all single-layer fully connected layers followed by Softmax. The semantic encoder $E_{s}$ is divided into two cases: When the semantics are text generated by LLM, $E_{s}$ is sentence-Bert, with initial parameters using all-MiniLM-L6-v2, and only the last two layers are fine-tuned during training. When the semantics are manually annotated attributes, $E_{s}$ is a two-layer fully connected layer with an output dimension of 15. The generator and discriminator of WGAN-GP are also both two-layer fully connected layers. The activation function of all the above-mentioned fully connected layers is ReLU. For the optimizers and various hyperparameters used in training, the specific details are as follows: $\lambda=1$, $T=5$, $\alpha=0.1$, $\beta=1$, the optimizer is Adam, and the learning rate is $2\times 10^{-4}$.
@@ IV-C Results on Office-home and Mini-DomainNet
We have adopted six comparative methods to compare with the proposed MDASR. Specifically, these six methods can be divided into CDZSL methods (including CuMix[22] and its two variants (img-only and two-level), S-AGG [35], and SEIC [24]) and generative method modified to apply in various domains (f-CLSWGAN [18]). CDZSL methods encompass the current approaches that aim to identify unseen classes in unseen domains. However, due to the influence of the DSP, they are not very effective at recognizing both seen and unseen classes simultaneously. Comparatively, generative methods can better overcome the impact of DSP, but it is not designed for CDGZSL. To ensure fairness, for comparative generative methods, we first utilize domain confusion to obtain a common space for multiple seen domains, and then generate features within the common space. Since our proposed method is a generative approach with meta semantic refinement, we also employed PCA[36] as another semantic refinement method for comparison.
@@
Methods Clipart Painting Sketch
@@
For the Mini-DomainNet dataset, due to the larger number of classes, we only demonstrate the results using semantics automatically acquired from LLMs, which are presented in Table III. Given that Mini-DomainNet has a greater number of unseen classes, the difficulty of CDGZSL is also higher. The proposed MDASR achieves an H-accuracy ranging from 36.62% to 40.25% in different unseen domains, which remains the highest among all the methods compared. The results of the other methods are similar to those on Office-Home but with a general decline in H-accuracy due to the difficulty of Mini-DomainNet. The DSP effect is still severe with CDZSL methods. Under the CDZSL setting, their accuracy for unseen classes may reach 8.50%-25.90% [24], but under the CDGZSL setting, it’s only about 2%-5%. f-CLSWGAN still maintains a certain capacity for alleviating DSP. If our MDASR is leveraged as the semantic refinement method, the H-accuracy is further improved compared to f-CLSWGAN; however, if PCA is used for semantic refinement, since its objective function is to maximize variance, which is unrelated to the intrinsic semantics, it usually brings about a negative effect.
Overall, in the two datasets, our proposed MDASR achieved the highest H-accuracy, presenting the capacity to recognize seen and unseen classes in unseen domains.
@@
Loss Clipart Painting Sketch
@@
Case (a): 61.38 24.53 35.05 58.23 23.31 33.29 56.60 22.03 31.71
Case (b): 61.34 29.32 39.67 58.29 20.23 30.03 57.32 26.02 35.79
Case (c): 63.73 26.15 37.08 59.98 23.82 34.10 55.41 26.25 35.62

### v2

@@ I Introduction
The era of Large Language Models (LLMs) marks a tremendous leap in the ability of artificial intelligence to understand and generate natural language. With massive numbers of parameters and extensive training data (for example, GPT-3 has 175 billion parameters and 753.4 GB of training data [1]), LLMs have the capacity to internalize vast amounts of knowledge and externalize it for application in the real world [2]. In some machine learning studies that fuse data and knowledge [3, 4, 5], researchers previously relied on the knowledge from human experts to provide models with prior information, aiming to obtain results that outperform purely data-driven techniques. The emergence of LLMs has, to some extent, supplanted the role of human experts, thereby bringing promising prospects to data-knowledge fusion studies, such as zero-shot learning (ZSL) [6].
ZSL is designed to recognize unseen classes not present in the training set by leveraging semantic knowledge to bridge the gap between seen and unseen classes, mimicking human cognitive inference [7, 8, 9]. Initially, ZSL approaches like DAP [8] and ALE [10] construct mappings from data to semantics, training on seen classes and extrapolating to unseen test data, under the assumption that only unseen classes would be tested. This works well when the assumption holds, but real-world scenarios often involve a mix of seen and unseen classes. To overcome this limitation, Generalized Zero-Shot Learning (GZSL) [11, 12, 13, 14] is introduced, requiring models to identify both seen and unseen classes during testing. However, ZSL models tend to misclassify unseen classes as seen due to the domain shift problem (DSP) [15, 16], leading to poor GZSL performance. To combat this, generative methods such as f-CLSWGAN [12] and ABAGAN [17] have been developed, modeling between the semantic space and the feature space, generating unseen class data based on semantic knowledge to reduce model overfitting on seen classes and mitigate DSP.
@@
ZSL Single Unseen class only
GZSL Single Seen & unseen class
CDZSL Multiple Unseen class only
CDGZSL Multiple Seen & unseen class
@@ III-F The CDGZSL inference procedure
Because $C_{c}$ is trained by both real seen features and fake unseen features with less effection of semantic redundancy, the DSP could be further alleviated, and we could gain a more accurate classifier of CDGZSL.
@@ IV-A2 Mini-DomainNet
{tblr}
colspec=c, cell11 = r=2, cell12 = c=2,r=2, cell14 = c=3, cell17 = c=3, cell110 = c=3, cell31 = r=8, cell32 = r=3, cell62 = c=2, cell72 = c=2, cell82 = r=2, cell111 = r=8, cell112 = r=3, cell142 = c=2, cell152 = c=2, cell162 = r=2, cell182 = c=2, vline2 = 1-18, hline1,3,11,19 = -, hline2 = 4-12, hline4-5,9,13,19-20 = 3, hline6-8,10,11-12,14-15,16-18,21-22 = 2-3, Semantics & Methods Art Clipart Product
@@
LLM text CuMix img-only 59.43 0 0 50.91 3.28 6.16 66.21 0.72 1.42
two-level 57.37 1.85 3.58 50.77 3.61 6.74 66.88 0.70 1.38
original 57.83 2.46 4.71 51.19 3.94 7.31 65.09 1.80 3.50
@@
Attribute CuMix img-only 43.81 4.93 8.86 41.10 2.96 5.52 56.49 2.52 4.82
two-level 43.45 5.56 9.85 39.35 4.60 8.23 55.17 6.49 11.61
original 44.28 6.17 10.83 38.95 7.23 12.19 53.16 9.38 15.94
@@ IV-B Experimental details
In CDGZSL, there are both seen and unseen domains, as well as seen and unseen classes. For Office-Home, the allocation of seen and unseen domains is as follows: Art, Clipart, and Product are each designated as the unseen domain in turn. When a domain is designated as the unseen domain, the other three domains serve as the seen domains. Because we employ Resnet-50 pretrained on Imagenet[33] as the initial parameters for the data encoder $E_{d}$, only the Real-world domain is not used as the unseen domain (since it is more similar to Imagenet) among these four domains. Additionally, the division of seen and unseen classes is as follows: we select 5 classes that do not appear in Imagenet as unseen classes, 5 classes as validation, and the remaining 55 classes as seen classes. For Mini-DomainNet, similarly, aside from the Real domain, the other three domains each take a turn as the unseen domain once. Also, we select 13 classes that do not appear in Imagenet as unseen classes, 13 classes as validation, and the remaining classes as seen classes.
For the various networks used in the experiments, the specific details are as follows: The data encoder $E_{d}$ is the encoder of Resnet50, with initial parameters trained on Imagenet; The domain classifier $C_{d}$, seen class classifier $C_{s}$, and the final CDGZSL classifier $C_{c}$ are all single-layer fully connected layers followed by Softmax. The semantic encoder $E_{s}$ is divided into two cases: When the semantics are text generated by LLM, $E_{s}$ is sentence-Bert [26], with initial parameters using all-MiniLM-L6-v2, and only the last two layers are fine-tuned during training. When the semantics are manually annotated attributes, $E_{s}$ is a two-layer fully connected layer with an output dimension of 15. The generator and discriminator of WGAN-GP are also both two-layer fully connected layers. The activation function of all the above-mentioned fully connected layers is ReLU. For the optimizers and various hyperparameters used in training, the specific details are as follows: $\lambda=1$, $T=5$, $\alpha=0.1$, $\beta=1$, the optimizer is Adam, and the learning rate is $2\times 10^{-4}$.
@@ IV-C Results on Office-home and Mini-DomainNet
We have adopted six comparative methods to compare with the proposed MDASR. Specifically, these six methods can be divided into CDZSL methods (including CuMix[20] and its two variants (img-only and two-level), S-AGG [34], and SEIC [22]) and generative method modified to apply in various domains (f-CLSWGAN [12]). CDZSL methods encompass the current approaches that aim to identify unseen classes in unseen domains. However, due to the influence of the DSP, they are not very effective at recognizing both seen and unseen classes simultaneously. Comparatively, generative methods can better overcome the impact of DSP, but it is not designed for CDGZSL. Since our proposed method is a generative approach with meta semantic refinement, we also employed PCA[35] as another semantic refinement method for comparison.
@@
{tblr}
cell11 = c=2,r=2, cell13 = c=3, cell16 = c=3, cell19 = c=3, cell31 = r=3, cell61 = c=2, cell71 = c=2, cell81 = r=2, cell111 = c=2, hline1,3,11 = -, hline2 = 3-11, hline4-5,9-10 = 2, hline6-8,10-11 = 1-2, Methods & Clipart Painting Sketch
@@
For the Mini-DomainNet dataset, due to the larger number of classes, we only demonstrate the results using semantics automatically acquired from LLMs, which are presented in Table III. Given that Mini-DomainNet has a greater number of unseen classes, the difficulty of CDGZSL is also higher. The proposed MDASR achieves an H-accuracy ranging from 36.62% to 40.25% in different unseen domains, which remains the highest among all the methods compared. The results of the other methods are similar to those on Office-Home but with a general decline in H-accuracy. The DSP effect is still severe with CDZSL methods. Under the CDZSL setting, their accuracy for unseen classes may reach 8.50%-25.90% [22], but under the CDGZSL setting, it’s only about 2%-5%. f-CLSWGAN still maintains a certain capacity for alleviating DSP. If our MDASR is leveraged, the H-accuracy is further improved compared to f-CLSWGAN; however, if PCA is used for semantic refinement, since its objective function is to maximize variance, which is unrelated to the intrinsic semantics, it usually brings about a negative effect. Overall, in the two datasets, our proposed MDASR achieved the highest H-accuracy, presenting the capacity to recognize seen and unseen classes in unseen domains.
@@
{tblr}
cell11 = c=3, cell14 = c=3, cell17 = c=3, cell110 = c=3, hline1,3,7 = -, hline2 = 2-13, & Loss Clipart Painting Sketch
@@
Case (a): ✘ ✘ 61.38 24.53 35.05 58.23 23.31 33.29 56.60 22.03 31.71
Case (b): ✘ 61.34 29.32 39.67 58.29 20.23 30.03 57.32 26.02 35.79
Case (c): ✘ 63.73 26.15 37.08 59.98 23.82 34.10 55.41 26.25 35.62